# Quantum Chemical and Trajectory Surface Hopping Molecular Dynamics Study of Iodine-based BODIPY Photosensitizer

### v1

@@ Abstract
A computational study of I-BODIPY (2-ethyl-4,4-difluoro-6,7-diiodo-1,3-dimethyl-4-bora-3a,4a-diaza-s-indacene) has been carried out to investigate its key photophysical properties as a potential triplet photosensitizer capable of generating singlet oxygen. Multireference CASPT2 and CASSCF methods have been used to calculate vertical excitation energies and spin–orbit couplings (SOCs), respectively, in a model (monoiodinated BODIPY) molecule to assess the applicability of the single-reference second-order algebraic diagrammatic construction, ADC(2), method to this and similar molecules. Subsequently, time-dependent density functional theory (TD-DFT), possibly within the Tamm–Dancoff approximation (TDA), using several exchange-correlation functionals has been tested on I-BODIPY against ADC(2), both employing a basis set with a two-component pseudopotential on the iodine atoms. Finally, the magnitudes of SOC between excited electronic states of all types found have thoroughly been discussed using the Slater–Condon rules applied to an arbitrary one-electron one-center effective spin–orbit Hamiltonian. The geometry dependence of SOCs between the lowest-lying states has also been addressed. Based on these investigations, the TD-DFT/B3LYP and TD-DFT(TDA)/BHLYP approaches have been selected as the methods of choice for the subsequent absorption spectra and trajectory surface hopping (TSH) molecular dynamics (MD) simulations, respectively. Two bright states in the visible spectrum of I-BODIPY have been found, exhibiting a red shift of the main peak with respect to unsubstituted BODIPY caused by the iodine substituents. Excited-state MD simulations including both nonadiabatic effects and SOCs have been performed to investigate the relaxation processes in I-BODIPY after its photoexcitation to the S1 state. The TSH MD simulations revealed that intersystem crossings occur on a time scale comparable to internal conversions and that after an initial phase of triplet population growth a “saturation” is reached where the ratio of the net triplet to singlet populations is about 4:1. The calculated triplet quantum yield of 0.85 is in qualitative agreement with the previously reported experimental singlet oxygen generation yield of 0.99$\pm$0.06.
@@ 2.1 Methods and basis sets
Turbomole has been used for geometry optimization of I-BODIPY in the ground state S0 by DFT with the B3LYP functional employing the dhf-TZVP basis set. Gaussian 09 48 has been used for geometry optimization and harmonic vibrational analysis of I-BODIPY (also in the ground state S0 by DFT but mainly) in the excited states S1 and T1 by TD-DFT with the B3LYP and M06-2X functionals due to the availability of analytical second derivatives. The aug-cc-pVDZ basis set 49, 50 has been employed for all atoms except iodine, for which we used the aug-cc-pVDZ-PP basis set 51, 47. For all geometry optimizations, the Grimme’s dispersion correction was taken into account. 52.
ADC(2) calculations have been carried out by the program Turbomole employing the dhf-TZVP basis set. DKH2 (second-order Douglas–Kroll–Hess) CASSCF and CASPT2 calculations have been performed using the program Molcas 7 53 employing the TZP contracted ANO-RCC basis set. 54 DMRG calculations have been done using the program MOLMPS55 employing the dhf-TZVP basis set. The program BDF 56, 57 has been used mainly for the calculation of electron densities in the ground and the lowest-excited singlet states by the sf-X2C-S-TD-DFT method 58 with the B3LYP functional employing the x2c-TZVPPall basis set. 59
For absorption spectra simulations and FSSH MD simulations including ISCs 60, 61 the program Newton-X 62, 63 with our implementation of nonadiabatic dynamics with time-derivative couplings 64 and spin–orbit couplings 65 applying the 3-step integrator approach 66 has been used. All simulations have been performed using the dhf-TZVP basis set.
In TSH MD, often decoherence corrections are employed 67, 68, 69. However, they were designed for localized, quickly damped couplings and their use might be problematic when SOCs are included 70, 71. We thus did not use any decoherence correction in our simulations.
@@ 2.2 Spin-orbit couplings
The employed effective one-electron spin–orbit Hamiltonian is the sum of the Breit–Pauli or DKH1 (first-order Douglas–Kroll–Hess) one-electron spin–orbit Hamiltonian 72 (which itself is a sum of contributions of individual atoms of a molecule) and a basis set independent approximation to the “two-electron” part of the ‘exact’ Breit–Pauli or DKH1 one-electron mean-field spin–orbit Hamiltonian 73. The approximation is a gentle modification of the so-called flexible nuclear screening spin–orbit (FNSSO) approximation 74 showing mean relative error below a few percent for virtually all elements of the periodic table in most bonding situations. A more descriptive view at the approximation is based on the explicit form of e.g. the Breit–Pauli one-electron spin–orbit Hamiltonian 72 (in SI units)
@@
The contribution of a heavy atom on which a two-component (relativistic, $j$-dependent) pseudopotential 75 is defined (and a reduced number of basis functions are centered to represent just the pseudovalence orbitals instead of the full set of core and valence orbitals) to matrix elements of the effective one-electron spin–orbit Hamiltonian is best evaluated using an operator implicitly included in the two-component pseudopotential itself. 76
The central idea behind the pseudopotential approximation 77, 78 is to select from the total number of $n$ (actually indistinguishable!) electrons of a molecule the $q$ valence electrons (responsible for all the chemistry) that are effectively moving in the electric field of the remaining $n-q$ core electrons and the positively charged nuclei (to which the core electrons are “stuck” in exactly the same way as in the isolated atoms). Heavy nuclei are thus replaced with spherical cores, while $Q_{C}e$ is the charge of the core whose center is at the nucleus $C$ with the charge $Z_{C}e$, $0<Q_{C}<Z_{C}$. Consequently,
@@
The pseudopotential $\Delta\mathrm{v}_{C}$ is a function of the distance of an electron from the nucleus $C$ and is generally different for basis functions of different angular symmetries. It is repulsive in the short range (so as to keep the valence electrons out of the core) and attractive in the long range. Typically, $Q_{C}$ differs from $Z_{C}$ only for the atoms of the heaviest elements, often just one or two atoms in the molecule. For atoms with $Q_{C}=Z_{C}$ the pseudopotential $\Delta\mathrm{v}_{C}$ clearly vanishes and a full one-electron basis is required. The contribution of such an atom to matrix elements of the effective one-electron spin–orbit Hamiltonian is then calculated using the (Breit–Pauli variant of the) one-center FNSSO approximation described above. For atoms with $Q_{C}<Z_{C}$ the related part of the one-center effective spin–orbit Hamiltonian is extracted from the two-component pseudopotential $\Delta\mathrm{v}_{C}$ by spin separation 76.
@@ 2.3 Accelerated nonadiabatic MD approach to ISC
Given that iodinated BODIPY derivatives are known to exhibit ISCs on the picosecond timescale, 28 we employed an accelerated approach in our simulations. Following the works of K. A. Lingerfelt et al. 79 and A. Nijamudheen et al. 80, we assume that the rates of slow nonadiabatic transitions are reasonably well captured by the Fermi’s golden rule which describes the transition rate $k$ as being proportional to the square of the matrix element $V$ of the related perturbation operator between the initial and final adiabatic states,
@@
This approach allows the acceleration of slow relaxation processes to a time window (of about 1 ps) more practical for NAMD simulations. In Refs. 79, 80 the scaling was applied to nonadiabatic couplings (NACs). However, in our case the slow process is ISC governed by SOC while the same-spin nonadiabatic transitions are fast enough and need no acceleration. We have thus applied the scaling to SOCs only, leaving NACs unchanged (as it was done also in our previous works 39, 38).
@@ 3.2 Benchmarking TD-DFT functionals
M. R. Momeni et al. 81 calculated the vertical excitation energies of the lowest singlet S1 in 17 BODIPY derivatives both by TD-DFT using 9 density functionals and by a plethora of ab initio methods including TD-HF, CIS, CIS(D), EOM-CCSD, SAC-CI, CC2, LR-CCSD, CCSDR(T), CASSCF or CASPT2 (the last two within several active spaces up to 12 electrons in 11 orbitals) emploing the cc-pVDZ and cc-pVTZ basis sets and compared them with experimental values. The observed failure of TD-DFT (mean absolute error greater than 0.3 eV) was attributed to the detected different amounts of electron correlation in the ground and excited states as well as to the finding that, while the ground state could in all cases be well treated by single-reference coupled cluster methods, the excited state S1 had, quite naturally and hardly surprisingly, a multireference character (i.e., was represented by a multiconfiguration wave function) and, in some cases, might not be described accurately enough without including double excitations from the dominant ground-state configuration.
M. de Vetta et al. 82 evaluated the performance of ADC(2) on the unsubstituted BODIPY molecule employing the aug-cc-pVDZ basis set against multistate CASPT2 using the active space of 12 electrons in 11 orbitals and the TZ contracted ANO-L basis set and found that the position of the dominant transition in the spectrum obtained by CASPT2 was 0.18 eV blue-shifted from the experimental value while ADC(2) yielded a further 0.13 eV blue shift with respect to CASPT2. Notice, though, that CASPT2 using a small active space and basis lacking diffuse functions can yield a result inferior to ADC(2). 83.
In a recent paper, V. Postils et al. 84 chose basically all the methods and density functionals used in the three papers mentioned above and, employing the cc-pVTZ basis set, calculated the vertical excitation energies of the electronic states S1 and T1 of unsubstituted BODIPY to uncover why not only TD-DFT but also some wave-function-based methods cannot accurately predict their energies. The main reason was reported to be a mild open-shell character of (i.e., strong HOMO and LUMO exchange interaction in) the ground state of the molecule.
@@
As an additional test of the possible multireference character of the ground states of BODIPY derivatives, we performed a DMRG calculation of the ground state of monoiodinated BODIPY in the position 2 using (frozen) RHF canonical orbitals and an active space consisting of 32 electrons in 30 orbitals with bond dimension 512. The resulting orbital entropies 85 were all below 0.3, and only for HOMO and LUMO their values exceeded 0.2, which can be seen in Table S25. The leading CI expansion coefficient (reconstructed from the MPS wave function) was 0.95 while the next largest one (in absolute value) of $-0.12$ belonged to the double excitation from HOMO to LUMO. For I-BODIPY we carried out an analogous DMRG calculation (48e in 46o, bond dimension 1024) and the resulting orbital entropies gave a qualitatively identical picture (cf. Table S26). We thus see no clear reason why single-reference methods like ADC(2) or TD-DFT should be rejected a priori for methods applicable to I-BODIPY and similar molecules.
@@
Vertical excitation energies of the low-lying excited states of monoiodinated BODIPY in the position 2 calculated by DKH2 CAS(18,15)PT2 are given in Table S1a. Symmetries and characters (of the dominant excitations) of the excited electronic states are included as well. For comparison, similar calculations have been done also by the aforementioned less demanding single-reference (i.e., black-box) methods ADC(2) and TD-DFT employing the B3LYP, BHLYP, and M06-2X functionals. The results of these calculations are given in Tables S1b through S1g. Besides the excitation energy and oscilator strength the so-called orbital overlap has also been calculated (by numerical integration on a grid) for each excited state as the sum of the overlap integrals of the absolute values of a pair of MOs (the one from which an electron was promoted and the one to which the electron was promoted) weighted by the square of the corresponding CIS expansion coefficient of the particular auxiliary wave function. 86 The orbital overlaps, whose values lie between 0 and 1, are measures of locality of the (single) excitations dominating the ADC(2) and TD-DFT excited states. Their small values, typically less than 0.3, may indicate either a charge-transfer (CT) or a Rydberg state.
@@
Vertical excitation energies, oscilator strengths, and orbital overlaps of the low-lying excited states of I-BODIPY at the optimized S0, S1 and T2 geometries calculated by ADC(2) and by TD-DFT using the B3LYP, BHLYP and M06-2X functionals are summarized in Tables S2a through S8c. Since there are two iodine substituents in I-BODIPY, there are also twice as many molecular orbitals located on iodine atoms (see Table S12 with the shapes and energies of the frontier MOs calculated by KS SCF employing the B3LYP functional), twice as many $(n,\pi^{*})$ states, more than twice as many $(\pi,\sigma^{*})$ states, and four times as many $(n,\sigma^{*})$ states compared to monoiodinated BODIPY. The number of $(\pi,\pi^{*})$ states is also greater due to the additional AO 5p of iodine perpendicular to the plane of the rings. However, not all of these “new” states have always been found in the selected energy windows. Nevertheless, slightly higher density of excited states is observed in I-BODIPY compared to monoiodinated BODIPY in the position 2.
@@
In all but a few calculations (carried out exclusively by TD-DFT employing the B3LYP functional), for both the BODIPY derivatives, the lowest three excited singlets and the lowest four triplets are $(\pi,\pi^{*})$ states. The order of the higher excited states depends, often dramatically, on the method (and density functional) used. The dependence is most pronounced for the $(n,\pi^{*})$ states whose lowest calculated triplet excitation energy in the monoiodinated BODIPY molecule varies from 3.51 eV obtained by TD-DFT/B3LYP for T5 to 4.70 eV obtained, coincidentally, by both ADC(2) and TD-DFT/BHLYP (or 4.59 eV obtained by TD-DFT/M06-2X) for T6 to 4.95 eV obtained by CAS(18,15)PT2 for T7. It is nevertheless fair to say that the last value may be slightly overestimated due to the fact that the doubly occupied MOs have been fixed at the values optimal for the lowest $(\pi,\sigma^{*})$ triplet (see above). Such a huge discrepancy, which also occurs for the $(n,\pi^{*})$ singlet as well as for the four lowest $(n,\pi^{*})$ states of I-BODIPY, can well be explained by the fact that all these states are dominated by excitations in the course of which an electron is promoted from an AO 5p of iodine perpendicular to the C–I bond and parallel to the plane of the rings to a distant $\pi^{*}$ orbital located above and below the rings (see also Table S12, namely HOMO-5, HOMO-3, and LUMO). And since it has long been known that linear response TD-DFT fails for CT states when used with “standard” exchange-correlation functionals like SVWN, BLYP or B3LYP, 87, 88 it is no surprise that the variability in the calculated excitation energies is the greatest precisely for the $(n,\pi^{*})$ states. This qualitative observation is in agreement with the calculated orbital overlaps whose values range between 0.45 and 0.75 for the vast majority of $(\pi,\pi^{*})$ states, between 0.5 and 0.6 for the $(n,\sigma^{*})$ states, and from 0.35 to 0.6 for the $(\pi,\sigma^{*})$ states (and a few $(\sigma,\pi^{*})$ states), but equal only about 0.15 for the lowest $(n,\pi^{*})$ state and are slightly above 0.2 for the second lowest $(n,\pi^{*})$ state (in the case of I-BODIPY) of each spin multiplicity.
@@
To avoid the failure of TD-DFT, i.e., a significant underestimation of the excitation energies of CT states, it is usually suggested that either long-range corrected hybrid functionals or hybrid functionals comprising about 50% of exact exchange (like the BHLYP and M06-2X functionals also included in our comparison) be used. 89, 90, 91, 92, 93 Indeed, the order of the lowest eight to ten excited singlets and triplets together with their excitation energies calculated by TD-DFT employing the M06-2X functional and by TD-DFT within the Tamm-Dancoff approximation (TDA) 94 employing the BHLYP functional are in best agreement with those calculated by ADC(2). Nevertheless, if the focus is on the three lowest excited singlets S1 through S3 of I-BODIPY at the ground-state optimized geometry, all dominated by $(\pi,\pi^{*})$ excitations, the energies calculated by TD-DFT with the B3LYP functional are the closest to the ADC(2) values.
@@
Vertical excitation energies and oscillator strengths of the lowest-lying singlets of I-BODIPY at its ground-state optimized geometry computed by ADC(2) and by TD-DFT with the B3LYP and BHLYP functionals have been summarized in Table 1. The ADC(2) vertical excitation energy 2.56 eV of the brightest state S1 is quite close to the experimental value of 2.46 eV. 9 Tables 1 and S2a through S7c compare the vertical excitation energies obtained by TD-DFT using various functionals with those calculated by ADC(2). Among the functionals tested, B3LYP gives the best agreement with ADC(2) for the vertical excitation energies of singlets S1 through S3, all dominated by $(\pi,\pi^{*})$ excitations. For singlets S4 through S7 dominated by either $(\pi,\sigma^{*})$ or $(n,\pi^{*})$ excitations (rarely in the same order), the energies calculated with the BHLYP and M06-2X functionals are the closest to the ADC(2) values while the energies calculated with the B3LYP functional are fatally underestimated. For the low-lying triplets, clearly the best agreement with the vertical excitation energies calculated by ADC(2) are observed for TD-DFT employing the M06-2X functional, closely followed by TD-DFT within TDA employing the BHLYP functional.

### v2

@@ Abstract
A computational study of I-BODIPY (2-ethyl-4,4-difluoro-6,7-diiodo-1,3-dimethyl-4-bora-3a,4a-diaza-s-indacene) has been carried out to investigate its key photophysical properties as a potential triplet photosensitizer capable of generating singlet oxygen. Multireference CASPT2 and CASSCF methods have been used to calculate vertical excitation energies and spin–orbit couplings (SOCs), respectively, in a model (monoiodinated BODIPY) molecule to assess the applicability of the single-reference second-order algebraic diagrammatic construction, ADC(2), method to this and similar molecules. Subsequently, time-dependent density functional theory (TD-DFT), possibly within the Tamm–Dancoff approximation (TDA), using several exchange-correlation functionals has been tested on I-BODIPY against ADC(2), both employing a basis set with a two-component pseudopotential on the iodine atoms. Finally, the magnitudes of SOC between excited electronic states of all types found have thoroughly been discussed using the Slater–Condon rules applied to an arbitrary one-electron one-center effective spin–orbit Hamiltonian. The geometry dependence of SOCs between the lowest-lying states has also been addressed. Based on these investigations, the TD-DFT/B3LYP and TD-DFT(TDA)/BHLYP approaches have been selected as the methods of choice for the subsequent nuclear ensemble approach absorption spectra simulations and mixed quantum-classical trajectory surface hopping (TSH) molecular dynamics (MD) simulations, respectively. Two bright states in the visible spectrum of I-BODIPY have been found, exhibiting a red shift of the main peak with respect to unsubstituted BODIPY caused by the iodine substituents. Excited-state MD simulations including both nonadiabatic effects and SOCs have been performed to investigate the relaxation processes in I-BODIPY after its photoexcitation to the S1 state. The TSH MD simulations revealed that intersystem crossings occur on a time scale comparable to internal conversions and that after an initial phase of triplet population growth a “saturation” is reached where the ratio of the net triplet to singlet populations is about 4:1. The calculated triplet quantum yield of 0.85 is in qualitative agreement with the previously reported experimental singlet oxygen generation yield of 0.99$\pm$0.06.
@@ 2.1 Methods and basis sets
Turbomole has been used for geometry optimization of both monoiodinated BODIPY in the position 2 and I-BODIPY in the ground state S0 by DFT with the B3LYP functional employing the dhf-TZVP basis set. Gaussian 09 48 has been used for geometry optimization of I-BODIPY in the excited states S1 and T2 by TD-DFT with the B3LYP functional due to the availability of analytical second derivatives. The aug-cc-pVDZ basis set 49, 50 has been employed for all atoms except iodine, for which we used the aug-cc-pVDZ-PP basis set 51, 47. For all geometry optimizations, the Grimme’s dispersion correction was taken into account. 52.
ADC(2) calculations have been carried out by the program Turbomole employing the dhf-TZVP basis set. DKH2 (second-order Douglas–Kroll–Hess) CASSCF and CASPT2 calculations have been performed using the program Molcas 7 53 employing the TZP contracted ANO-RCC basis set. 54 DMRG calculations have been done using the program MOLMPS55 employing the dhf-TZVP basis set. The program BDF 56, 57 has been used mainly for the calculation of electron densities in the ground and lowest-excited singlet states by the sf-X2C-S-TD-DFT method 58 with the B3LYP functional employing the x2c-TZVPPall basis set. 59
For nuclear ensemble approach (NEA) absorption spectra simulations60 and mixed quantum-classical Tully’s FSSH MD simulations61, 62, 63, 64 including ISCs 65, 66 the program Newton-X 67, 68 with our implementation of nonadiabatic dynamics with time-derivative couplings 69 and spin–orbit couplings 70 applying the 3-step integrator approach 71 has been used. All simulations have been performed using the dhf-TZVP basis set.
In TSH MD, often decoherence corrections are employed 72, 73, 74. However, they were designed for localized, quickly damped couplings and their use might be problematic when SOCs are included 75, 76. We thus did not use any decoherence correction in our simulations.
@@ 2.2 Spin-orbit couplings
The employed effective one-electron spin–orbit Hamiltonian is the sum of the Breit–Pauli or DKH1 (first-order Douglas–Kroll–Hess) one-electron spin–orbit Hamiltonian 77 (which itself is a sum of contributions of individual atoms of a molecule) and a basis set independent approximation to the “two-electron” part of the ‘exact’ Breit–Pauli or DKH1 one-electron mean-field spin–orbit Hamiltonian. 78 The approximation is a gentle modification of the so-called flexible nuclear screening spin–orbit (FNSSO) approximation 79 showing mean relative error below a few percent for virtually all elements of the periodic table in most bonding situations. A more descriptive view at the approximation is based on the explicit form of e.g. the Breit–Pauli one-electron spin–orbit Hamiltonian 77 (in SI units)
@@
The contribution of a heavy atom on which a two-component (relativistic, $j$-dependent) pseudopotential 80 is defined (and a reduced number of basis functions are centered to represent just the pseudovalence orbitals instead of the full set of core and valence orbitals) to matrix elements of the effective one-electron spin–orbit Hamiltonian is best evaluated using an operator implicitly included in the two-component pseudopotential itself. 81
The central idea behind the pseudopotential approximation 82, 83 is to select from the total number of $n$ (actually indistinguishable!) electrons of a molecule the $q$ valence electrons (responsible for all the chemistry) that are effectively moving in the electric field of the remaining $n-q$ core electrons and the positively charged nuclei (to which the core electrons are “stuck” in exactly the same way as in the isolated atoms). Heavy nuclei are thus replaced with spherical cores, while $Q_{C}e$ is the charge of the core whose center is at the nucleus $C$ with the charge $Z_{C}e$, $0<Q_{C}<Z_{C}$. Consequently,
@@
The pseudopotential $\Delta\mathrm{v}_{C}$ is a function of the distance of an electron from the nucleus $C$ and is generally different for basis functions of different angular symmetries. It is repulsive in the short range (so as to keep the valence electrons out of the core) and attractive in the long range. Typically, $Q_{C}$ differs from $Z_{C}$ only for the atoms of the heaviest elements, often just one or two atoms in the molecule. For atoms with $Q_{C}=Z_{C}$ the pseudopotential $\Delta\mathrm{v}_{C}$ clearly vanishes and a full one-electron basis is required. The contribution of such an atom to matrix elements of the effective one-electron spin–orbit Hamiltonian is then calculated using the (Breit–Pauli variant of the) one-center FNSSO approximation described above. For atoms with $Q_{C}<Z_{C}$ the related part of the one-center effective spin–orbit Hamiltonian is extracted from the two-component pseudopotential $\Delta\mathrm{v}_{C}$ by spin separation. 81
Adoption of the one-center approximation enabled us, first, to avoid implementing the existing, rather complicated, algorithm for the evaluation of multicenter matrix elements of the effective spin–orbit Hamiltonian implicitly included in a two-component pseudopotential between redundant Cartesian Gaussian functions 84 and, second, as a nice bonus, to exploit spherical symmetry. The derivation of the new, truly simple, formulae for the calculation of matrix elements of the one-electron one-center effective spin–orbit Hamiltonian implicitly included in a two-component pseudopotential between nonredundant Cartesian Gaussian functions is given in SI on pp. 110–115.
@@ 2.3 Accelerated nonadiabatic MD approach to ISC
Given that iodinated BODIPY derivatives are known to exhibit ISCs on the picosecond timescale, 28 we employed an accelerated approach in our simulations. Following the works of K. A. Lingerfelt et al. 85 and A. Nijamudheen et al. 86, we assume that the rates of slow nonadiabatic transitions are reasonably well captured by the Fermi’s golden rule which describes the transition rate $k$ as being proportional to the square of the matrix element $V$ of the related perturbation operator between the initial and final adiabatic states,
@@
This approach allows the acceleration of slow relaxation processes to a time window (of about 1 ps) more practical for NAMD simulations. In Refs. 85, 86 the scaling was applied to nonadiabatic couplings (NACs). However, in our case the slow process is ISC governed by SOC while the same-spin nonadiabatic transitions are fast enough and need no acceleration. We have thus applied the scaling to SOCs only, leaving NACs unchanged (as it was done also in our previous works 39, 38).
@@ 3.2 Benchmarking TD-DFT functionals
M. R. Momeni et al. 87 calculated the vertical excitation energies of the lowest singlet S1 in 17 BODIPY derivatives both by TD-DFT using 9 density functionals and by a plethora of ab initio methods including TD-HF, CIS, CIS(D), EOM-CCSD, SAC-CI, CC2, LR-CCSD, CCSDR(T), CASSCF or CASPT2 (the last two within several active spaces up to 12 electrons in 11 orbitals) emploing the cc-pVDZ and cc-pVTZ basis sets and compared them with experimental values. The observed failure of TD-DFT (mean absolute error greater than 0.3 eV) was attributed to the detected different amounts of electron correlation in the ground and excited states as well as to the finding that, while the ground state could in all cases be well treated by single-reference coupled cluster methods, the excited state S1 had, quite naturally and hardly surprisingly, a multireference character (i.e., was represented by a multiconfiguration wave function) and, in some cases, might not be described accurately enough without including double excitations from the dominant ground-state configuration.
M. de Vetta et al. 88 evaluated the performance of ADC(2) on the unsubstituted BODIPY molecule employing the aug-cc-pVDZ basis set against multistate CASPT2 using the active space of 12 electrons in 11 orbitals and the TZ contracted ANO-L basis set and found that the position of the dominant transition in the spectrum obtained by CASPT2 was 0.18 eV blue-shifted from the experimental value while ADC(2) yielded a further 0.13 eV blue shift with respect to CASPT2. Notice, though, that CASPT2 using a small active space and basis lacking diffuse functions can yield a result inferior to ADC(2). 89.
In a recent paper, V. Postils et al. 90 chose basically all the methods and density functionals used in the three papers mentioned above and, employing the cc-pVTZ basis set, calculated the vertical excitation energies of the electronic states S1 and T1 of unsubstituted BODIPY to uncover why not only TD-DFT but also some wave-function-based methods cannot accurately predict their energies. The main reason was reported to be a mild open-shell character of (i.e., strong HOMO and LUMO exchange interaction in) the ground state of the molecule.
@@
As an additional test of the possible multireference character of the ground states of BODIPY derivatives, we performed a DMRG calculation of the ground state of monoiodinated BODIPY in the position 2 using (frozen) RHF canonical orbitals and an active space consisting of 32 electrons in 30 orbitals with bond dimension 512. The resulting orbital entropies 91 were all below 0.3, and only for HOMO and LUMO their values exceeded 0.2, which can be seen in Table S25. The leading CI expansion coefficient (reconstructed from the MPS wave function) was 0.95 while the next largest one (in absolute value) of $-0.12$ belonged to the double excitation from HOMO to LUMO. For I-BODIPY we carried out an analogous DMRG calculation (48e in 46o, bond dimension 1024) and the resulting orbital entropies gave a qualitatively identical picture (cf. Table S26). We thus see no clear reason why single-reference methods like ADC(2) or TD-DFT should be rejected a priori for methods applicable to I-BODIPY and similar molecules.
@@
Vertical excitation energies of the low-lying excited states of monoiodinated BODIPY in the position 2 calculated by DKH2 CAS(18,15)PT2 are given in Table S1a. Symmetries and characters (of the dominant excitations) of the excited electronic states are included as well. For comparison, similar calculations have been done also by the aforementioned less demanding single-reference (i.e., black-box) methods ADC(2) and TD-DFT employing the B3LYP, BHLYP, and M06-2X functionals. The results of these calculations are given in Tables S1b through S1g. Besides the excitation energy and oscilator strength the so-called orbital overlap has also been calculated (by numerical integration on a grid) for each excited state as the sum of the overlap integrals of the absolute values of a pair of MOs (the one from which an electron was promoted and the one to which the electron was promoted) weighted by the square of the corresponding CIS expansion coefficient of the particular auxiliary wave function. 92 The orbital overlaps, whose values lie between 0 and 1, are measures of locality of the (single) excitations dominating the ADC(2) and TD-DFT excited states. Their small values, typically less than 0.3, may indicate either a charge-transfer (CT) or a Rydberg state.
@@
Vertical excitation energies, oscilator strengths, and orbital overlaps of the low-lying excited states of I-BODIPY at the optimized S0, S1 and T2 geometries calculated by ADC(2) and by TD-DFT using the B3LYP, BHLYP and M06-2X functionals are summarized in Tables S2a through S8c. Since there are two iodine substituents in I-BODIPY, there are also twice as many molecular orbitals located on iodine atoms see Figs. 2 through 10 and Table S12 with the shapes and energies of the frontier MOs calculated by KS SCF employing the B3LYP functional), twice as many $(n,\pi^{*})$ states, more than twice as many $(\pi,\sigma^{*})$ states, and four times as many $(n,\sigma^{*})$ states compared to monoiodinated BODIPY. The number of $(\pi,\pi^{*})$ states is also greater due to the additional AO 5p of iodine perpendicular to the plane of the rings. However, not all of these “new” states have always been found in the selected energy windows. Nevertheless, slightly higher density of excited states is observed in I-BODIPY compared to monoiodinated BODIPY in the position 2.
Figure 2: In-plane nonbonding ($n$) HOMO-5 of I-BODIPY, orbital energy -7.67 eV.
Figure 3: Out-of-plane nonbonding/$\pi$ bonding ($\pi$) HOMO-4 of I-BODIPY, orbital energy -7.65 eV.
Figure 4: In-plane nonbonding ($n$) HOMO-3 of I-BODIPY, orbital energy -7.20 eV.
Figure 5: $\pi$ bonding ($\pi$) HOMO-2 of I-BODIPY, orbital energy -7.07 eV.
Figure 6: Out-of-plane nonbonding/$\pi$ bonding ($\pi$) HOMO-1 of I-BODIPY, orbital energy -6.53 eV.
Figure 7: $\pi$ bonding ($\pi$) HOMO of I-BODIPY, orbital energy -5.98 eV.
Figure 8: $\pi$ antibonding ($\pi^{*}$) LUMO of I-BODIPY, orbital energy -3.07 eV.
Figure 9: C–I $\sigma$ antibonding ($\sigma^{*}$) LUMO+1 of I-BODIPY, orbital energy -1.38 eV.
Figure 10: C–I $\sigma$ antibonding ($\sigma^{*}$) LUMO+2 of I-BODIPY, orbital energy -0.55 eV.
@@
In all but a few calculations (carried out exclusively by TD-DFT employing the B3LYP functional), for both the BODIPY derivatives, the lowest three excited singlets and the lowest four triplets are $(\pi,\pi^{*})$ states. The order of the higher excited states depends, often dramatically, on the method (and density functional) used. The dependence is most pronounced for the $(n,\pi^{*})$ states whose lowest calculated triplet excitation energy in the monoiodinated BODIPY molecule varies from 3.51 eV obtained by TD-DFT/B3LYP for T5 to 4.70 eV obtained, coincidentally, by both ADC(2) and TD-DFT/BHLYP (or 4.59 eV obtained by TD-DFT/M06-2X) for T6 to 4.95 eV obtained by CAS(18,15)PT2 for T7. It is nevertheless fair to say that the last value may be slightly overestimated due to the fact that the doubly occupied MOs have been fixed at the values optimal for the lowest $(\pi,\sigma^{*})$ triplet (see above). Such a huge discrepancy, which also occurs for the $(n,\pi^{*})$ singlet as well as for the four lowest $(n,\pi^{*})$ states of I-BODIPY, can well be explained by the fact that all these states are dominated by excitations in the course of which an electron is promoted from an AO 5p of iodine perpendicular to the C–I bond and parallel to the plane of the rings to a distant $\pi^{*}$ orbital located above and below the rings (see also Table S12, namely HOMO-5, HOMO-3, and LUMO). And since it has long been known that linear response TD-DFT fails for CT states when used with “standard” exchange-correlation functionals like SVWN, BLYP or B3LYP, 93, 94 it is no surprise that the variability in the calculated excitation energies is the greatest precisely for the $(n,\pi^{*})$ states. This qualitative observation is in agreement with the calculated orbital overlaps whose values range between 0.45 and 0.75 for the vast majority of $(\pi,\pi^{*})$ states, between 0.5 and 0.6 for the $(n,\sigma^{*})$ states, and from 0.35 to 0.6 for the $(\pi,\sigma^{*})$ states (and a few $(\sigma,\pi^{*})$ states), but equal only about 0.15 for the lowest $(n,\pi^{*})$ state and are slightly above 0.2 for the second lowest $(n,\pi^{*})$ state (in the case of I-BODIPY) of each spin multiplicity.
@@
To avoid the failure of TD-DFT, i.e., a significant underestimation of the excitation energies of CT states, it is usually suggested that either long-range corrected hybrid functionals or hybrid functionals comprising about 50% of exact exchange (like the BHLYP and M06-2X functionals also included in our comparison) be used. 95, 96, 97, 98, 99 Indeed, the order of the lowest eight to ten excited singlets and triplets together with their excitation energies calculated by TD-DFT employing the M06-2X functional and by TD-DFT within the Tamm-Dancoff approximation (TDA) 100 employing the BHLYP functional are in best agreement with those calculated by ADC(2). Nevertheless, if the focus is on the three lowest excited singlets S1 through S3 of I-BODIPY at the ground-state optimized geometry, all dominated by $(\pi,\pi^{*})$ excitations, the energies calculated by TD-DFT with the B3LYP functional are the closest to the ADC(2) values.
@@
Vertical excitation energies and oscillator strengths of the lowest-lying singlets of I-BODIPY at its ground-state optimized geometry computed by ADC(2) and by TD-DFT with the B3LYP and BHLYP functionals have been summarized in Table 1. The ADC(2) vertical excitation energy 2.56 eV of the brightest state S1 is quite close to the experimental value of 2.46 eV. 9 Tables 1 and S2a through S7c compare the vertical excitation energies obtained by TD-DFT using various functionals with those calculated by ADC(2). Among the functionals tested, B3LYP gives the best agreement with ADC(2) for the vertical excitation energies of singlets S1 through S3, all dominated by $(\pi,\pi^{*})$ excitations. For singlets S4 through S7 dominated by either $(\pi,\sigma^{*})$ or $(n,\pi^{*})$ excitations (rarely in the same order), the energies calculated with the BHLYP and M06-2X functionals are the closest to the ADC(2) values while the energies calculated with the B3LYP functional are fatally underestimated. For the low-lying triplets, clearly the best agreement with the vertical excitation energies calculated by ADC(2) are observed for TD-DFT employing the M06-2X functional, closely followed by TD-DFT within TDA employing the BHLYP functional. However, it must be admitted that the agreement between ADC(2) and TD-DFT is never really good.